Consider the partial decomposition of a as 2A 2B + C. At equilibrium 700mL gaseous mixture contains 100mL of gas C at 10atm and 300K. What is the value if Kp for the reaction? A, B and C are all gases.
 
Can u please give a detailed explanation as i am struggling for 2 hrs on this.Thanks

Grade:10

1 Answers

Vikas TU
7 years ago
Deal it in either in Volume terms or Pressure terms.
Hence,
             2A ---------->  2B   +   C
t=0         a                 0          0
t=t        a -x               x         2x
at eqm. it is given,     C is 100 ml.
hence, 2x =100 ml.
 
x = 50 ml.
and also given,
a + 2x = 700 
a = 600 ml.
 
Kc = 50*100^2/450^2 = > 2.47 spprox.
Kp = Kc*(RT)^1 {delta n = 2 -1 }
Hence,
Kp = 2.47 * 0.0821 * 300
    = 61.4148.